# Uniform Field in Microwave Cavities Through the Use of Effective Magnetic Walls

### v3

@@ Abstract
Wire media (WM) resonators have emerged as promising realization for plasma haloscopes – devices designed to detect axions, a potential component of dark matter. Key factors influencing the detection probability include cavity volume, resonance quality factor, and form factor. While the form factor has been explored for resonant frequency tuning, its optimization for axion detection remains unexplored. In this work, we present a novel approach to significantly enhance the form factor of WM plasma haloscopes. By shifting the metal walls of the resonator by a quarter wavelength, we effectively convert an electric wall boundary condition into a magnetic wall one, allowing for an almost uniform mode. Theoretical analysis and numerical simulations confirm that this modification improves the electric field profile and boosts the form factor. We validate these findings through experimental results from two prototype resonators: one with a standard geometry and another with a quarter-wave air gap between the WM and the walls. Additionally, our method provides a simple way to control the field profile within WM cavities, which can be explored for further applications.
@@ I Introduction
Previous investigations into WM open resonators have primarily focused on exploiting topological transitions in their isofrequency contours to achieve a notable Purcell factor [9, 10]. More recently, cavities consisting in WM enclosed by metallic walls and placed in a DC magnetic field, often referred to as plasma haloscopes [11], have emerged as a promising technique for the detection of axions, a possible candidate for dark matter [12]. One can also look for dark photons, another promising dark matter candidate even if the magnetic field is absent [13]. Our earlier studies of WM-filled cavities demonstrated that WM operating in the epsilon-near-zero (ENZ) regime significantly extends the wavelength of light within the medium, resulting in a low-frequency fundamental resonance and a high degree of field uniformity [14].
@@
As to the axion haloscopes, the field uniformity is just as beneficial for them as for EPR devices. Several key characteristics should be engineered to enhance their performance: the strength of the DC magnetic field, cavity volume, quality factor, and form factor are critical factors influencing the power from photon-axion conversion in resonance [19]. The form factor for a mode in a plasma haloscope, characterized by an electric field $\mathbf{E}_{lmn}$, is defined as
[…]
where $V$ is the cavity volume, $\epsilon$ is the relative permittivity within the cavity, and $\mathbf{B}_{0}$ is the external DC magnetic field, which we will take to be spatially constant and aligned with the wires.
@@
In this work we show that placing the walls with the quarter-wavelength gap from the WM interface, we effectively transform this interface from the perfect electric conductor (PEC) into a perfect magnetic conductor (PMC). Figure 1 illustrates this concept. This transformation is similar to the conversion of a shortened transmission line into an open one by shifting the termination by an odd number of quarter-wavelengths. By modifying the boundary conditions, we create a cavity whose fundamental mode is $\mathrm{TM}_{000}$ rather than the $\mathrm{TM}_{110}$ mode of the regular cavity. This transition significantly improves the form factor and shifts the fundamental resonance frequency down to the plasma frequency.
@@ II.1 Effective boundary condition
To assess the impact of the air gap between the wall and the WM on the modal field uniformity, we develop a one-dimensional analytical model. We focus on the TM modes because the TE modes weakly interact with the WM. We divide the system into two regions, as depicted in Fig. 2. Region 1 contains the lossless (PEC wires) WM characterized by a relative permittivity [8]
$\epsilon=1-\left(\tfrac{k_{p}}{k_{0}}\right)^{2},$ (2)
where $k_{0}$ is the vacuum wave number and $k_{p}=\tfrac{2\pi}{\lambda_{p}}$, with $\lambda_{p}$ being the plasma wavelength. Region 2 represents the air gap between the WM and the PEC walls. Horizontal lines in Fig. 2 outline a unit cell of the infinitely extended parallel-plate cavity. By imposing the inherent odd symmetry of the electric field and even symmetry of the magnetic field due to the PECs, we derive the TM mode solutions to Maxwell’s equations as follows:
$E_{z}(x)=\begin{cases}A\cos{(k_{x1}x)}&|x|\leq\frac{d}{2}\\[5.0pt] B\sin{\left(k_{2x}\left(\tfrac{d}{2}+g-|x|\right)\right)}&|x|>\frac{d}{2}\end{cases},$ (3)
$H_{y}(x)=\begin{cases}-jA\frac{k_{1x}}{\eta_{0}k_{0}}\sin{(k_{x1}x)}&|x|\leq\frac{d}{2}\\[5.0pt] -j\frac{x}{|x|}B\frac{k_{2x}}{\eta_{0}k_{0}}\cos{\left(k_{x2}\left(\tfrac{d}{2}+g-|x|\right)\right)}&|x|>\frac{d}{2}\end{cases},$ (4)
where $k_{x1}=\sqrt{\epsilon}k_{0}$, and $k_{x2}=k_{0}$, assuming no field dependence on $x$ and $z$. We evaluate the field solutions in Eqs. (3) and (4) at $x=\tfrac{d}{2}^{+}$ (on the side of the air gap) to derive a surface impedance that represents the effective boundary condition for the WM interface,
$Z=\frac{|E_{z}|(\tfrac{d}{2}^{+})}{|H_{y}|(\tfrac{d}{2}^{+})}=\frac{\eta_{0}k_{0}}{k_{x2}}\frac{\sin{(k_{x2}g)}}{\cos{(k_{x2}g)}}.$ (5)
@@ II.2 Fundamental mode of WM resonator
[…]
@@ II.3 Results
We used full-wave simulations in a COMSOL Multiphysics eigenmode solver to validate the applicability of our 1D theory to a 2D cavity. In both our analytical calculations and simulations, the cavity comprised an array of 10$\times$10 infinitely long wires with wire radius $r_{0}$ = 0.5 mm and inter-wire spacing $a$ = 10 mm. We adopted the generalized quasi-static approach from Ref. [27] to determine the requisite plasma frequency for the theoretical model.
@@
Figure 5 illustrates the resonant frequency (Fig. 5a) and form factor (Fig. 5b) of the fundamental mode in such WM resonator as a function of the air gap, $g$. The resonant frequency remains nearly constant for $40\leq g\leq\lambda_{p}/$ 4, aligning with the plasma frequency at $g=\lambda_{p}/4$. Conversely, the form factor progressively increases, reaching a maximum at an air gap of approximately $g=\lambda_{p}/4$. While the resonant frequency exhibits excellent agreement between analytical and numerical results, coming from both a physical WM and an effective WM, within this range, a slight discrepancy arises in the form factor. This error results from the evident difference between the effective and physical (real) WM. In the effective-medium model there are no wires and the field does not vanish inside this medium, whereas in the physical WM it vanishes within every wire – see the dark spots in the bottom panel of Fig. 3b. In accordance with numerical calculations of the mode field within a single unit cell of our WM the maximally achievable form factor corresponding at 8.9 GHz to $g=\lambda_{p}/4=$8.4 mm is equal to 0.92, as well as in the analytical model. However, the simulation of the physical WM corresponding to the bottom panel of Fig. 3b shows the form factor 0.89. Despite this, the maximum form factor occurs at nearly the same air gap value in both numerical simulations and the analytical model. For smaller gaps the agreement between the analytical and numerical results corresponding to physical WM is good, and is excellent for the numerical simulations of the effective WM sample surrounded by the PEC walls. Thus, these results fully justify the model of an effective WM sample for $g\leq\lambda_{p}/4$. For $g>\lambda_{p}/4$, the agreement between the analytical model and numerical simulations disappears due to the effect of the field concentration in the air gap. This effect invalidates the initial assumption of field confinement within the WM sample. However, the agreement between the results obtained for the effective WM sample and for the array of separate wires keeps.
@@ III Experiment
We equipped each resonator prototype with a 4-mm long SMA connector attached to its bottom plate for signal feeding. The SMA port functions similarly to a monopole antenna, exciting the TM modes within the resonator cavity. To characterize the electric field distribution of the TM modes, we positioned a separate monopole antenna through a pre-drilled array of holes on the top plate of each resonator, as illustrated Fig. 6b. This monopole antenna was a coaxial cable with a center conductor radius of 0.25 mm, extending 4 mm into the resonators through the holes. The hole array comprised 81 individual holes, arranged in a square grid with a period of $a=1$ cm and a hole radius of $r_{h}=3$ mm.
Figure 7: Experimentally measured $\mathrm{S}_{21}$ parameter for wire media resonators: (a) Regular resonator and (b) $\lambda_{p}/4.2$-shifted resonator. Measurements were performed by placing the movable monopole antenna in the central hole. The resonance frequency of the fundamental TM mode in the regular resonator ($f_{110}$) is $9.11$ GHz (experiment) and $9.10$ GHz (simulation), while in the $\lambda_{p}/4.2$-shifted resonator ($f_{000}$) it is $8.943$ GHz (experiment and simulation).
We measured the S-parameters using a Rohde &\&& Schwarz ZVB20 vector network analyzer (VNA) across a frequency range of 8.75-9.75 GHz. The fixed antenna, located on the bottom plate of each prototype, was connected to one port of the VNA, while the movable antenna on the top plate was connected to the other port. To measure the $\mathrm{S_{21}}$, we positioned the monopole antenna in the array of holes. Fig. 7 displays the measured $\mathrm{S_{21}}$ when the movable antenna was placed in the central hole of each prototype. We observe that the spectrum of the $\lambda_{p}/4.2$-shifted resonator is red-shifted relative to that of the regular resonator, as predicted by theory for the fundamental mode. Meanwhile, the amplitude levels and shape of the spectrum change slightly, indicating that the quality factor of the resonator is not significantly affected when an air gap of $\lambda_{p}/4.2$ is introduced.
@@
We characterized the operation of the two resonators under comparison by examining the first peaks of the $\mathrm{S_{21}}$-parameter spectra. These peaks correspond to the $\mathrm{TM_{110}}$ mode for the regular resonator and the quasi-$\mathrm{TM_{000}}$ mode for the $\lambda_{p}/4.2$-shifted resonator, as shown in Fig. 7. We do not consider high-order modes, as they are not relevant for our purpose of achieving the highest field homogeneity (see e.g. in Refs. [17, 14]). Fig. 8 presents the measured electric field distributions for both prototypes at the corresponding frequencies ($f_{000}$ for the $\lambda_{p}/4.2$-shifted resonator and $f_{110}$ for the regular one). We observe a significant improvement in field homogeneity for the $\lambda_{p}/4.2$-shifted resonator compared to the regular resonator. This observation is further supported by the measured form factors, with a value of $0.77$ (experiment), $0.80$ (simulation) for the $\lambda_{p}/4.2$-shifted resonator compared to $0.61$ (experiment), $0.64$ (simulation) for the regular resonator.
@@ IV Conclusions
We introduced a new method to enhance field homogeneity within WM resonators. Our theoretical model, supported by numerical simulations, demonstrated that properly increasing the air gap between the resonator walls and the WM sample, one can effectively manipulate the field homogeneity. When the air gap reached one-quarter of the resonant wavelength, the field homogeneity peaked, theoretically resulting in a form factor of 0.92 for an effective WM model (theory and simulations) and 0.89 for a physical array of wires (simulations). Additionally, this enhancement in the form factor was accompanied by a shift in the resonant mode frequency toward the plasma frequency.

### v4

@@ Abstract
Wire media (WM) resonators have emerged as promising realization for plasma haloscopes – devices designed to detect axions, a potential component of dark matter. Key factors influencing the detection probability include cavity volume, resonance quality factor, and form factor. While the form factor has been explored for resonant frequency tuning, its optimization for axion detection remains unexplored. In this work, we present a novel approach to significantly enhance the form factor of WM plasma haloscopes. By shifting the metal walls of the resonator by a quarter wavelength, we effectively convert an electric wall boundary condition into a magnetic wall one, allowing for an almost uniform mode. Theoretical analysis and numerical simulations confirm that this modification improves the electric field profile and boosts the form factor, while also slightly enhancing the quality factor. We validate these findings through experimental results from two prototype resonators: one with a standard geometry and another with a quarter-wave air gap between the WM and the walls. Additionally, our method provides a simple way to control the field profile within WM cavities, which can be explored for further applications.
@@ I Introduction
Previous investigations into WM open resonators have primarily focused on exploiting topological transitions in their isofrequency contours to achieve a notable Purcell factor [9, 10]. More recently, cavities consisting in WM enclosed by metallic walls and placed in a DC magnetic field, often referred to as plasma haloscopes [11], have emerged as a promising technique for the detection of axions, a possible candidate for dark matter [12]. One can also look for dark photons, another promising dark matter candidate, even if the magnetic field is absent [13]. Additionally, the detection of high-frequency gravitational waves can be explored [14, 15]. Our earlier studies of WM-filled cavities demonstrated that WM operating in the epsilon-near-zero (ENZ) regime significantly extends the wavelength of light within the medium, resulting in a low-frequency fundamental resonance and a high degree of field uniformity [16].
@@
When the axion frequency matches the resonant frequency of a microwave cavity mode, the power generated by axion-photon conversion is given by [21, 12]:
$P_{a}\approx g_{a\gamma\gamma}^{2}\frac{\rho_{a}}{m_{a}}B_{0}^{2}VC_{lmn}\mathrm{min}(Q_{L},Q_{a})$ (1)
where $g_{a\gamma\gamma}$ is the axion-photon coupling constant, $m_{a}$ is the axion mass, $\rho_{a}$ is the local mass density of the axion field, and $\mathbf{B}_{0}$ is the external DC magnetic field. The parameters $V$ and $Q_{L}$ denote the cavity volume and loaded quality factor, respectively, while $Q_{a}$ represents the quality factor of the galactic halo axion signal. $C_{lmn}$ is the form factor associated with the $\mathrm{TM}_{lmn}$ mode of the cavity. To maximize the axion-photon conversion power in axion haloscopes, it is crucial to optimize the external DC magnetic field, cavity volume, quality factor, and form factor. Notably, the scanning rate of an axion search experiment, which probes a range of axion masses, is proportional to the square of the form factor [22].
For axion haloscopes, field uniformity is as crucial as it is for EPR devices, since the form factor directly quantifies this uniformity. The form factor for a mode in a plasma haloscope, characterized by an electric field $\mathbf{E}_{lmn}$, is defined as
$C_{lmn}=\frac{\left|\iiint_{V}\mathbf{B}_{0}\cdot\mathbf{E}_{lmn}\,dV\right|^{2}}{VB_{0}^{2}\iiint_{V}{\epsilon|\mathbf{E}_{lmn}|^{2}}\,dV},$ (2)
where $\epsilon$ is the relative permittivity within the cavity, and $\mathbf{B}_{0}$ is assumed to be spatially constant and aligned with the wires.
@@
In this work we show that placing the walls with the quarter-wavelength gap from the WM interface, we effectively transform this interface from the perfect electric conductor (PEC) into a perfect magnetic conductor (PMC). Figure 1 illustrates this concept. This transformation is similar to the conversion of a shortened transmission line into an open one by shifting the termination by an odd number of quarter-wavelengths. By modifying the boundary conditions, we create a cavity whose fundamental mode is $\mathrm{TM}_{000}$ rather than the $\mathrm{TM}_{110}$ mode of the regular cavity. This transition significantly improves the form factor and shifts the fundamental resonance frequency down to the plasma frequency. Here, we define $\mathrm{TM}_{lmn}$ modes as those with an electric field parallel to the wire axes and a magnetic field perpendicular to them.
@@ II.1 Effective boundary condition
To assess the impact of the air gap between the wall and the WM on the modal field uniformity, we develop a one-dimensional analytical model. We consider the WM as an effective uniaxial medium, whose permittivity tensor is given by [8]:
$\hat{\epsilon}=\begin{pmatrix}\epsilon_{t}&0&0\\ 0&\epsilon_{t}&0\\ 0&0&\epsilon_{z}\end{pmatrix}$ (3)
We focus on TM modes, since TE modes weakly interact with the WM when the wire radius, $r$, is much smaller than the lattice period, $a$, i.e., $r\ll a$, for which $\epsilon_{t}\approx 1$. The TM modes have an electric field parallel to the wire axes ($z$-axis) and a magnetic field perpendicular to them ($y$-axis). We divide the system into two regions, as depicted in Fig. 2. Region 1 contains the lossless (PEC wires) WM characterized by a relative permittivity [8]
$\epsilon\equiv\epsilon_{z}|_{k_{z}=0}=1-\left(\tfrac{k_{p}}{k_{0}}\right)^{2},$ (4)
where $k_{0}$ is the vacuum wave number, $k_{z}$ is the $z$-component of the wave vector, and $k_{p}=\tfrac{2\pi}{\lambda_{p}}$, with $\lambda_{p}$ being the plasma wavelength. Region 2 represents the air gap between the WM and the PEC walls. Horizontal lines in Fig. 2 outline a unit cell of the infinitely extended parallel-plate cavity. By imposing the inherent odd symmetry of the electric field and even symmetry of the magnetic field due to the PECs, we derive the TM mode solutions to Maxwell’s equations as follows:
$E_{z}(x)=\begin{cases}A\cos{(k_{x1}x)}&|x|\leq\frac{d}{2}\\[5.0pt] B\sin{\left(k_{x2}\left(\tfrac{d}{2}+g-|x|\right)\right)}&|x|>\frac{d}{2}\end{cases},$ (5)
$H_{y}(x)=\begin{cases}-jA\frac{k_{x1}}{\eta_{0}k_{0}}\sin{(k_{x1}x)}&|x|\leq\frac{d}{2}\\[5.0pt] -j\frac{x}{|x|}B\frac{k_{x2}}{\eta_{0}k_{0}}\cos{\left(k_{x2}\left(\tfrac{d}{2}+g-|x|\right)\right)}&|x|>\frac{d}{2}\end{cases},$ (6)
where $k_{x1}=\sqrt{\epsilon}k_{0}$, and $k_{x2}=k_{0}$, assuming no field dependence on $y$ and $z$. We evaluate the field solutions in Eqs. (5) and (6) at $x=\tfrac{d}{2}^{+}$ (on the side of the air gap) to derive a surface impedance that represents the effective boundary condition for the WM interface,
$Z=\frac{E_{z}(\tfrac{d}{2}^{+})}{H_{y}(\tfrac{d}{2}^{+})}=\frac{j\eta_{0}k_{0}}{k_{x2}}\frac{\sin{(k_{x2}g)}}{\cos{(k_{x2}g)}}.$ (7)
@@ II.2 Fundamental mode of WM resonator
$C=\frac{\left(\iint_{S}{E_{z}}\,dS\right)^{2}}{S\iint_{S}{|\mathbf{E}|^{2}}\,dS},$ (13)
@@ II.3 Results
We used full-wave simulations in a COMSOL Multiphysics eigenmode solver to validate the applicability of our 1D theory to a 2D cavity. In both our analytical calculations and simulations, the cavity comprised an array of $\times10$ infinitely long wires with wire radius $r$ = 0.5 mm and inter-wire spacing $a$ = 10 mm. We adopted the generalized quasi-static approach from Ref. [29] to determine the requisite plasma frequency for the theoretical model, which expresses the plasma frequency as a function of $r$ and $a$.
@@
Figure 5 illustrates the resonant frequency (Fig. 5a) and form factor (Fig. 5b) of the fundamental mode in such WM resonator as a function of the air gap, $g$. The resonant frequency remains nearly constant for $0\leq g\leq\lambda_{p}/4$, aligning with the plasma frequency at $g=\lambda_{p}/4$. Conversely, the form factor progressively increases, reaching a maximum at an air gap of approximately $g=\lambda_{p}/4$. While the resonant frequency exhibits excellent agreement between analytical and numerical results, coming from both a physical WM and an effective WM, within this range, a slight discrepancy arises in the form factor. This error results from the evident difference between the effective and physical (real) WM. In the effective-medium model there are no wires and the field does not vanish inside this medium, whereas in the physical WM it vanishes within every wire – see the dark spots in the bottom panel of Fig. 3b.
In accordance with numerical calculations of the mode field of our WM resonator the maximally achievable form factor corresponding at 8.9 GHz to $g=\lambda_{p}/4=$8.4 mm is equal to 0.92, as well as in the analytical model. However, the simulation of the physical WM corresponding to the bottom panel of Fig. 3b shows the form factor 0.89. Despite this, the maximum form factor occurs at nearly the same air gap value in both numerical simulations and the analytical model. Our strategy considerably enhances the form factor of microwave cavities for axion detection, surpassing the maximum form factor of 0.69 found in cylindrical cavities [22, 24].
For smaller gaps the agreement between the analytical and numerical results corresponding to physical WM is good, and is excellent for the numerical simulations of the effective WM sample surrounded by the PEC walls. Thus, these results fully justify the model of an effective WM sample for $g\leq\lambda_{p}/4$. For $g>\lambda_{p}/4$, the agreement between the analytical model and numerical simulations disappears due to the effect of the field concentration in the air gap. This effect invalidates the initial assumption of field confinement within the WM sample. However, the agreement between the numerical results obtained for the effective WM sample and for the array of separate wires keeps.
@@
To analyze the quality factor, we focused solely on losses due to the wires, as they are the dominant loss mechanism in WM cavities [16]. We considered copper wires with a conductivity of $\sigma=5.7\times 10^{7}~\mathrm{S/m}$ for both the analytical model and the 2D numerical simulations. By using the formulas for the quality factor described in Ref. [16], along with the updated plasma frequency and wire inductance from Ref. [29], we obtained the analytical results. Introducing an air gap of a quarter-wavelength in the wire media resonator significantly enhances the form factor and slightly increases the quality factor, as shown in Fig. 6. When the air gap exceeds a quarter-wavelength, the quality factor rapidly increases as the field becomes primarily concentrated in the newly added air gap region.
Figure 6: Quality factor of wire medium (WM) resonator fundamental TM mode for different air gap thicknesses, $g$. Wires were modeled as copper with a conductivity of $\sigma=5.7\times 10^{7}~\mathrm{S/m}$.
@@ III Experiment
We equipped each resonator prototype with a 4-mm long SMA connector attached to its bottom plate for signal feeding. The SMA port functions similarly to a monopole antenna, exciting the TM modes within the resonator cavity. To characterize the electric field distribution of the TM modes, we positioned a separate monopole antenna through a pre-drilled array of holes on the top plate of each resonator, as illustrated Fig. 6b. This monopole antenna was a coaxial cable with a center conductor radius of 0.25 mm, extending 4 mm into the resonators through the holes. The hole array comprised 81 individual holes, arranged in a square grid with a period of $a=1$ cm and a hole radius of $r_{h}=3$ $m$.
Figure 8: Experimentally measured $\mathrm{S}_{21}$ parameter for wire media resonators: (a) Regular resonator and (b) $\lambda_{p}/4.2$-shifted resonator. Measurements were performed by placing the movable monopole antenna in the central hole. The resonance frequency of the fundamental TM mode in the regular resonator ($f_{110}$) is $9.11$ GHz (experiment) and $9.10$ GHz (simulation), while in the $\lambda_{p}/4.2$-shifted resonator ($f_{000}$) it is $8.943$ GHz (experiment and simulation). The insets show $|E_{z}|$ in the xyxy plane (as in Fig. 1) at $10~\mathrm{mm}$ above the center of the $\lambda_{p}/4.2$-shifted resonator. In the color maps, yellow represents 1 and dark blue represents 0 in arbitrary units (a.u.) for $|E_{z}|$.
We measured the S-parameters using a Rohde &\& Schwarz ZVB20 vector network analyzer (VNA) across a frequency range of 8.75-9.75 GHz. The fixed antenna, located on the bottom plate of each prototype, was connected to one port of the VNA, while the movable antenna on the top plate was connected to the other port. To measure the $\mathrm{S_{21}}$, we positioned the monopole antenna in the array of holes. Fig. 8 displays the measured $\mathrm{S_{21}}$ when the movable antenna was placed in the central hole of each prototype. We observe that the spectrum of the $\lambda_{p}/4.2$-shifted resonator is red-shifted relative to that of the regular resonator, as predicted by theory for the fundamental mode. In general, the $\mathrm{TM}_{lmn}$ modes in the spectrum of the $\lambda/4.2$-shifted resonator are red-shifted compared to those of the regular resonator, allowing the $\mathrm{TM}_{02n}^{c}$ modes to fall within the studied frequency range, as detailed in Appendix A. The superscript $c$ in the $\mathrm{TM}_{02n}^{c}$ modes indicates a cross-coupling between the $x$ and $y$ dependence of the fields, as discussed in Appendix A.
@@
We characterized the operation of the two resonators under comparison by examining the first peaks of the $\mathrm{S_{21}}$-parameter spectra. These peaks correspond to the $\mathrm{TM_{110}}$ mode for the regular resonator and the quasi-$\mathrm{TM_{000}}$ mode for the $\lambda_{p}/4.2$-shifted resonator, as shown in Fig. 8. We do not consider high-order modes, as they are not relevant for our purpose of achieving the highest field homogeneity (see e.g. in Refs. [19, 16]). Fig. 9 presents the measured electric field distributions for both prototypes at the corresponding frequencies ($f_{000}$ for the $\lambda_{p}/4.2$-shifted resonator and $f_{110}$ for the regular one). We observe a significant improvement in field homogeneity for the $\lambda_{p}/4.2$-shifted resonator compared to the regular resonator. This observation is further supported by the measured form factors, with a value of $0.77$ (experiment), $0.80$ (simulation) for the $\lambda_{p}/4.2$-shifted resonator compared to $0.61$ (experiment), $0.64$ (simulation) for the regular resonator. The regular resonator consists of a cavity filled exclusively with WM, as described in Ref. [16].
@@ IV Conclusions
We introduced a new method to enhance field homogeneity within WM resonators. Our theoretical model, supported by numerical simulations, demonstrated that properly increasing the air gap between the resonator walls and the WM sample, one can effectively manipulate the field homogeneity. When the air gap reached one-quarter of the resonant wavelength, the field homogeneity peaked, theoretically resulting in a form factor of 0.92 for an effective WM model (theory and simulations) and 0.89 for a physical array of wires (simulations). Additionally, this enhancement in the form factor was accompanied by a shift in the resonant mode frequency toward the plasma frequency, along with a slight increase in the quality factor.
@@
In the case of a dark matter search, the signal power is directly proportional to the form factor. If one achieves the maximum improvement of $40\%$ in the form factor the scanning rate of the experiment, which goes as the square of the power, is improved by almost a factor of two, especially when one considers the increased quality factor as well. This improvement comes at no real additional costs, unlike say increasing the size or strength of the magnetic field. Further, these improvements are not mutually exclusive with other enhancements such as higher magnetic fields or lower noise floors. Particularly at high frequencies improving the scanning speed of dark matter experiments is crucial to search for dark matter.